How do I solve the equation 4x^2 = 20. And how do I find the x-intercepts are after I solve the equation.
Answer by haileytucki(390)

4x^(2)=20 IN THIS EQUATION, THE / SIGN STANDS FOR +- AND THE ~ SIGN STANDS FOR THE SQUARE ROOT OF SYMBOL.
Divide each term in the equation by 4.
(4x^(2))/(4)=(20)/(4)
Simplify the left-hand side of the equation by canceling the common factors.
x^(2)=(20)/(4)
Simplify the right-hand side of the equation by simplifying each term.
x^(2)=5
Take the square root of both sides of the equation to eliminate the exponent on the left-hand side.
x=\~(5)
First, substitute in the + portion of the \ to find the first solution.
x=~(5)
Next, substitute in the - portion of the \ to find the second solution.
x=-~(5)
The complete solution is the result of both the + and - portions of the solution.
x=~(5),-~(5)_x=2.24,-2.24